For each scenario, calculate the income elasticity of demand, determine whether the good is inferior or normal, and classify the goods income elasticity. When ca answers to the nearest hundredth g the income elasticity of demand, use the midpoint formula. Round your Sylvias annual salary increases from $100,500 to $109,500, and she decides to increase the number of vacations she takes per year from three to four. Calculate her income elasticity of demand for vacations Vacations are a income elasticity yacations As a good, vacations are Blake eats two bags of generic potato chips each day, and does not purchase any name-brand chips. Blakes hourly wage increases from $8.75 to $15, and he decides to eat one name-brand bag and one generic-brand bag each day. Calculate Blakes income elasticity of demand for generic potato chips Generic potato chips are a income elasticity generic chips As a good, generic potato chips are
Answer #1

Answer 1:

➡️Income elasticity of demand for VACATION.

=(Q2-Q1) /[(Q2+Q1) /2] ÷ (Y2-Y1) /[(Y2+Y1) /2]

=(4-3)/[(4+3)/2] ÷(109500 - 100500)/[(109500+100500)/2]

=1/3.5 ÷ [9000/105000]

=0.29/0.086

=3.37

➡️VACATIONS ARE. . . .

Answer : elastic

Explanation : if income elasticity is more than 1, it is elastic...... But if it is less than 1 than it is inelastic.... So here the income elasticity is more than one.. SO IT IS ELASTIC.

➡️AS A GOOD VACATIONS ARE....

Answer : it is a normal good.

Explanation: if good is normal, by increase in income raises the quantity demanded. Becouse quantity demanded and income moves in the same direction .. So it is a normal good.

Answer 2:

➡️Income elasticity of demand for GENERIC CHIPS

=(Q2-Q1) /[(Q2+Q1) /2] ÷ (Y2-Y1) /[(Y2+Y1) /2]

=(1-2)/[(1+2)/2] ÷ (15-8.75)/[(15+8.75)/2]

= - 1/1.5 ÷ 6.25/11.875

=- 0.67/ 0.53

= - 1.26

➡️ Generic potato chips are....

Answer : Elastic

Explanation : becouse here the income elasticity of demand is greater than one.

➡️As a good generic potato chips are...

Answer : inferior good.

Explanation:

Here, by increase in income, the quantity demand for the generic potato chips falles, so it is the inferior good. In the case of inferior goods higher the income lowers the demanded. Because the quantity demanded of inferior goods and income move in opposite direction.